In the morning, Maria drove to a business appointment at 30 mph. Her average speed on the return trip in the afternoon was 20 mph. The return trip took 1/2 hour longer because of heavy traffic. How far did she travel to the appointment?
Please help me answer this question.

Answer by TimothyLamb(4379)

---
s = d/t
d = s*t
---
x = outbound time
---
outbound:
30x = d
---
inbound:
20(x + 0.5) = d
20x + 10 = d
---
30x = 20x + 10
10x = 10
x = 1 hour
---
d = 30x
---
answer:
d = 30 miles
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php